# Communication-Efficient and Privacy-Adaptable Mechanism for Federated Learning

### v1

@@ III CEPAM for FL
3) Server: The server observes $\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}}$ from client $k$. Using the shared randomness between client $k$ and the server, i.e., the shared random seed $s_{k}$, the same realizations of $\mathbf{V}_{H_{t+\tau,j}^{k}}$’s and $U_{t+\tau,j}^{k}$’s generated by client $k$ can also be obtained by the server. More specifically, the decoder generates $U_{t+\tau,j}^{k}\sim g$ and $\mathbf{V}_{t+\tau,j,1},\ldots,\mathbf{V}_{t+\tau,j,H_{t+\tau,j}^{k}}\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$ by the PRNG $\tilde{\mathfrak{P}}^{k}$, and outputs $\mathbf{Y}_{t+\tau,j}^{k}=\beta(U_{t+\tau,j}^{k})(\mathbf{M}_{t+\tau,j}^{k}+\mathbf{V}_{H_{t+\tau,j}^{k}})$. Subsequently, the decoder collects the sub-vectors $\{\mathbf{Y}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ into a $m\times 1$ and scales it by $\gamma^{-1}\|\mathbf{X}_{t+\tau}^{k}\|$ to obtain an estimation $\hat{\mathbf{X}}_{t+\tau}^{k}$ of the update $\mathbf{X}_{t+\tau}^{k}$. The decoding algorithm is summarized in Algorithm 3.
Algorithm 3 $\mathrm{DECODE}(\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}},\gamma,s_{k},g,N)$
1:Inputs: Set of messages $\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}}$, scaling coefficient $\gamma$, random seed $s_{k}$, pdf $g$, number of sub-vectors $N$
@@
3:Use $\tilde{\mathfrak{P}}^{k}$ to sample $U_{t+\tau,j}^{k}\sim g$triangleright Initiated by seed $s_{k}$
@@
6:     Use $\tilde{\mathfrak{P}}^{k}$ to sample $\mathbf{V}_{t+\tau,j,1},\ldots,\mathbf{V}_{t+\tau,j,H_{t+\tau,j}^{k}}\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$triangleright Initiated by seed $s_{k}$
7:      Compute $\mathbf{Y}_{t+\tau,j}^{k}\leftarrow\beta(U_{t+\tau,j}^{k})(\mathbf{M}_{t+\tau,j}^{k}+\mathbf{V}_{H_{t+\tau,j}^{k}})$triangleright Perform decoding, inducing privacy protection
@@
9:Collect $\{\mathbf{Y}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ into $\mathbf{Y}_{t+\tau}^{k}$
10:Estimate $\hat{\mathbf{X}}_{t+\tau}^{k}\leftarrow\gamma^{-1}\|\mathbf{X}_{t+\tau}^{k}\|\cdot\mathbf{Y}_{t+\tau}^{k}$
11:return $\hat{\mathbf{X}}_{t+\tau}^{k}$
Privacy enhancement: The privacy is ensured by LRSUQ executed cooperatively between the clients and the server in compliance with the threat model and the defined privacy guarantee. Through the decoding process, it is ensured that $\{\mathbf{Y}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ comprises noisy estimates of $\{\tilde{\mathbf{X}}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$, i.e., the estimated model updates at the server are noisy estimates of the model updates at the clients, thus establishing a privacy mechanism. Refer to Step 7 in Algorithm 3. Specifically, in Section IV-A1, we prove that when using the Gaussian mechanism, the global average of estimated model updates satisfies $(\epsilon,\delta)$-DP requirement by specifying appropriate $f$ and $g$, while in Section IV-A2, when using the Laplace mechanism, the estimated model updates satisfies $\epsilon$-DP requirement by specifying appropriate $f$ and $g$.
IV Performance Analysis
In this section, we study the performance of CEPAM. This includes characterizing its privacy guarantees and compression capabilities, followed by an exploration of its distortion bounds and convergence analysis. We begin by stating a useful lemma for the subsequent analysis. The proof is given in Appendix -A.
Lemma 11
@@ IV-A Privacy
IV-A1 Gaussian mechanism
If $g$ follows a chi-squared distribution with $n+2$ degrees of freedom, i.e., $g\sim\chi_{n+2}^{2}$, then $f$ follows a Gaussian distribution. This guarantees that the global average of estimated model updates satisfies $(\epsilon,\delta)$-DP. The proof is given in Appendix -B.
Theorem 12
Set $\tilde{U}_{t+\tau,j}^{k}\sim g=\chi_{n+2}^{2}$ and $\tilde{\mathbf{Z}}_{t+\tau,j}^{k}|\{\tilde{U}_{t+\tau,j}^{k}=u\}\sim\mathrm{Unif}(\sigma\sqrt{u}B^{n})$ for every $k,j$, the resulting mechanism, i.e., Step 7 in Algorithm 3, is Gaussian. The average $\frac{1}{K}\sum_{k\in\mathcal{K}}\hat{\mathbf{X}}_{t+\tau}^{k}$ is a noisy estimate of the average of the model updates such that
$\frac{1}{K}\sum_{k\in\mathcal{K}}\hat{\mathbf{X}}_{t+\tau}^{k}=\frac{1}{K}\sum_{k\in\mathcal{K}}\mathbf{X}_{t+\tau}^{k}+\mathcal{N}\bigg{(}\mathbf{0},\frac{\sigma^{2}\|\mathbf{X}_{t+\tau}^{k}\|^{2}}{K\gamma^{2}}\mathbf{I}_{m}\bigg{)},$ (15)
@@ IV-A1 Gaussian mechanism
IV-A2 Laplace mechanism
If $g$ follows a Gamma distribution $\mathrm{Gamma}(2,1)$, then $f$ follows a Laplace distribution.444 Our CEPAM can be extended to include multivariate Laplace distributions [59] or $t$-distributions [60], which are left for future work. This guarantees that estimated model updates satisfy $\epsilon$-DP. The proof is given in Appendix -C.
Theorem 13
Set $\tilde{U}_{t+\tau,j}^{k}\sim g=\mathrm{Gamma}(2,1)$ and $\tilde{Z}_{t+\tau,j}^{k}|\{\tilde{U}_{t+\tau,j}^{k}=u\}\sim\mathrm{Unif}((-bu,bu))$ for every $k$ and $j$, the resulting mechanism is Laplace. The estimator $\hat{\mathbf{X}}_{t+\tau}^{k}$ is a noisy estimate of the model update $\mathbf{X}_{t+\tau}^{k}$ such that
$\hat{\mathbf{X}}_{t+\tau}^{k}=\mathbf{X}_{t+\tau}^{k}+\mathrm{Lap}\left(\mathbf{0},(b\gamma^{-1}\|\mathbf{X}_{t+\tau}^{k}\|)\mathbf{I}_{m}\right),$ (17)
@@ IV-B Compression
IV-C Distortion bounds
Representing the clipped model update $\tilde{\mathbf{X}}_{t+\tau}^{k}$ of client $k$ using a finite number of bits inherently introduces distortion, i.e., the recovered vector $\hat{\mathbf{X}}_{t+\tau}^{k}=\tilde{\mathbf{X}}_{t+\tau}^{k}+\mathbf{Z}_{t+\tau}^{k}$. We can then characterize the squared norm of the FL quantization error. The proof is given in Appendix -D.
Proposition 14
The FL quantization error $\mathbf{Z}_{t+\tau}^{k}$ has zero mean and satisfies
$\mathbb{E}[\|\mathbf{Z}_{t+\tau}^{k}\|^{2}\big{|}\mathbf{x}_{t+\tau}^{k}]=\gamma^{-2}\|\mathbf{x}_{t+\tau}^{k}\|^{2}N\mathrm{Var}(f).$ (18)
At the end of each FL round, CEPAM incorporates a DP mechanism based on LRSUQ, inherently introducing distortion as mentioned above. This distortion emerges as the update $\mathbf{X}_{t+\tau}^{k}$ is mapped into its distorted counterpart $\hat{\mathbf{X}}_{t+\tau}^{k}$. Consequently, the global model (4) becomes
$\hat{\mathbf{W}}_{t+\tau}:=\mathbf{W}_{t}+\sum_{k\in\mathcal{K}}p_{k}\hat{\mathbf{X}}_{t+\tau}^{k}=\sum_{k\in\mathcal{K}}p_{k}\hat{\mathbf{W}}_{t+\tau}^{k},$ (19)
where $\hat{\mathbf{W}}_{t+\tau}^{k}=\mathbf{W}_{t}+\hat{\mathbf{X}}_{t+\tau}^{k}$. Under common assumptions used in FL analysis, we can bound the error between $\mathbf{W}_{t+\tau}$ and $\hat{\mathbf{W}}_{t+\tau}$. To that end, we adopt the following assumptions:
AS1: The $\xi_{k,j}$’s, where $j=1,\ldots,n_{k}$, are i.i.d. samples in $\mathcal{D}^{(k)}$. However, different datasets $\mathcal{D}^{(k)}$ can be statistically heterogeneous and follow different distributions.
AS2: Let the sample index $j_{t}^{k}$ be sampled from client $k$’s local data $\mathcal{D}^{(k)}$ uniformly at random. The expected squared norm of stochastic gradients is bounded by some $\theta_{k}^{2}>0$, i.e., $\mathbb{E}[\|\nabla F_{k}^{j_{t}^{k}}(\mathbf{W})\|]\leq\theta_{k}^{2}$, for all $\mathbf{W}\in\mathbb{R}^{m}$.
We have the following bound on the error between $\mathbf{W}_{t+\tau}$ and $\hat{\mathbf{W}}_{t+\tau}$. The proof is given in Appendix -E.
Proposition 15
When AS2 holds, the mean-squared error $\mathbb{E}[\|\hat{\mathbf{W}}_{t+\tau}-\mathbf{W}_{t+\tau}\|^{2}]$ is upper-bounded by
$N\gamma^{-2}\tau\mathrm{Var}(f)\sum_{t^{\prime}=t}^{t+\tau-1}\eta_{t^{\prime}}^{2}\sum_{k\in\mathcal{K}}p_{k}^{2}\theta_{k}^{2}.$ (20)
IV-D FL Convergence analysis
We next study the FL convergence of CEPAM, considering the following common assumptions used in the FL [40, 41, 9]:
AS3: $F_{1},\ldots,F_{K}$ are all $L$-smooth, i.e., for all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{m}$, $F_{k}(\mathbf{y})-F_{k}(\mathbf{x})\leq(\mathbf{y}-\mathbf{x})^{T}\nabla F_{k}(\mathbf{x})+\frac{L}{2}\|\mathbf{y}-\mathbf{x}\|^{2}$.
AS4: $F_{1},\ldots,F_{K}$ are all $C$-strongly convex, [^5] The extension to non-convex objective functions is left for future work. i.e., for all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{m}$, $F_{k}(\mathbf{y})-F_{k}(\mathbf{x})\geq(\mathbf{y}-\mathbf{x})^{T}\nabla F_{k}(\mathbf{x})+\frac{C}{2}\|\mathbf{y}-\mathbf{x}\|^{2}$.
We define the following to quantify the degree of non-iid:
$\psi=F(\mathbf{w}^{*})-\sum_{k\in\mathcal{K}}\min_{\mathbf{w}\in\mathbb{R}^{m}}F_{k}(\mathbf{w}),$ (21)
where $\mathbf{w}^{*}$ denotes the minimum of $F(\cdot)$. If the data is iid, meaning that the training data originates from the same distribution, then $\psi$ goes to zero as the training size increases. In the case of non-i.i.d. data, $\psi$ is positive, and its magnitude reflects the heterogeneity of the data distribution.
The convergence of CEPAM for FL learning is given as follows. The proof is given in Appendix -F.
Theorem 16
Assume that AS1-4 hold and $\theta_{k}$ $L$, $C$ be defined therein. Set $\alpha=\tau\max\{\frac{4L}{C},1\}$ and the learning rate $\eta_{t}=\frac{\zeta}{t+\alpha}$ for nonnegative integers $t\geq 0$. Then, CEPAM satisfies
$\mathbb{E}[F(\mathbf{W}_{T})]-F(\mathbf{w}^{*})\leq\frac{L}{2(T+\alpha)}\max\left\{\frac{B\tau^{2}}{C^{2}(\tau-1)},\alpha\|\mathbf{W}_{0}-\mathbf{w}^{*}\|\right\},$ (22)
where
$B:=6L\psi+\left(4N\gamma^{-2}\tau^{2}\mathrm{Var}(f)+1\right)\sum_{k\in\mathcal{K}}p_{k}^{2}\theta_{k}^{2}+8(\tau-1)^{2}\sum_{k\in\mathcal{K}}p_{k}\theta_{k}^{2}.$
V Numerical Evaluations
@@
In this section, we evaluate the performance of CEPAM. [^6] The source code used in our numerical evaluations is available at https://github.com/yokiwuuu/CEPAM.git. We begin by detailing our experimental setup, which includes the types of datasets, model architectures, and training configurations in Section V-A. Afterwards, we present comprehensive experimental results to demonstrate the effectiveness of CEPAM by comparing it to several baselines realizing the privacy protection and quantization in the FL framework in Section V-B.
@@
We also compare the accuracy-privacy trade-off between CEPAM and the Gaussian-mechanism-then-quantize approach in Section V-C. All experiments were executed on a server equipped with dual Intel(R) Xeon(R) Platinum 8480C CPUs (112 cores in total), 512GB RAM, and an NVIDIA RTX A5000 GPU (24GB VRAM) running Ubuntu 22.04.4 LTS. The implementation was based on PyTorch and CUDA 12.5 with NVIDIA driver version 555.42.06.
V-A Experimental Setup
V-A1 Datasets
@@ V-A3 Baselines
  • FL+{Gaussian, Laplace}: FL with one-dimensional Gaussian or Laplace mechanism, no compression
  • FL+{Gaussian, Laplace}+SDQ: FL with approach that applies one-dimensional Gaussian or Laplace mechanism followed by scalar SDQ.
  • CEPAM-{Gaussian, Laplace}: CEPAM achieves privacy and quantization jointly through LRSUQ to construct Gaussian or Laplace mechanisms, namely CEPAM-Gaussian or CEPAM-Laplace. In particular, for CEPAM-Gaussian, we evaluate three different cases of LRSUQ that simulate Gaussian noise for dimension $n=1,2,3$, where we use integer lattice $\mathbb{Z}^{n}$, [^7] To simplify implementation, we choose to use the integer lattice due to its well-established decoding algorithm [62]. The time complexity of the rejection sampling step can be improved by using a lattice with higher packing density. and the corresponding basic cells $(-0.5,0.5]$, $(-0.5,0.5]^{2}$, and $(-0.5,0.5]^{3}$. For CEPAM-Laplace using LRSUQ to simulate Laplace noise, we use integer lattice $\mathbb{Z}$ with the basic cell $(-0.5,0.5]$.
V-A4 Training Configurations

### v2

@@ IV CEPAM for FL
3) Server: The server observes $\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}}$ from client $k$. Using the shared randomness between client $k$ and the server, i.e., the shared random seed $s_{k}$, the same realizations of $\mathbf{V}_{H_{t+\tau,j}^{k}}$’s and $U_{t+\tau,j}^{k}$’s generated by client $k$ can also be obtained by the server. More specifically, the decoder generates $U_{t+\tau,j}^{k}\sim g$ and $\mathbf{V}_{t+\tau,j,1},\ldots,\mathbf{V}_{t+\tau,j,H_{t+\tau,j}^{k}}\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$ by the PRNG $\tilde{\mathfrak{P}}^{k}$, and outputs $\mathbf{Y}_{t+\tau,j}^{k}=\beta(U_{t+\tau,j}^{k})(\mathbf{M}_{t+\tau,j}^{k}+\mathbf{V}_{H_{t+\tau,j}^{k}})$. Subsequently, the decoder collects the sub-vectors $\{\mathbf{Y}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ into an estimated vector $\hat{\mathbf{X}}_{t+\tau}^{k}$. of the update $\mathbf{X}_{t+\tau}^{k}$. The decoding algorithm is summarized in Algorithm 3.
Algorithm 3 $\mathrm{DECODE}(\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}},s_{k},g,N)$
1:Inputs: Set of messages $\{(H_{t+\tau,j}^{k},\mathbf{M}_{t+\tau,j}^{k})\}_{j\in\mathcal{N}}$, random seed $s_{k}$, pdf $g$, number of sub-vectors NN
@@
3:Use $\tilde{\mathfrak{P}}^{k}$ to sample $U_{t+\tau,j}^{k}\sim g$ $\triangleright$ Initiated by seed $s_{k}$
@@
6:  Use $\tilde{\mathfrak{P}}^{k}$ to sample $\mathbf{V}_{t+\tau,j,1},\ldots,\mathbf{V}_{t+\tau,j,H_{t+\tau,j}^{k}}\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$ $\triangleright$ Initiated by seed $s_{k}$
7:   Compute $\mathbf{Y}_{t+\tau,j}^{k}\leftarrow\beta(U_{t+\tau,j}^{k})(\mathbf{M}_{t+\tau,j}^{k}+\mathbf{V}_{H_{t+\tau,j}^{k}})$ $\triangleright$ Perform decoding, inducing privacy protection
@@
9:Collect $\{\mathbf{Y}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ into $\mathbf{Y}_{t+\tau}^{k}$ and set $\hat{\mathbf{X}}_{t+\tau}^{k}\leftarrow\mathbf{Y}_{t+\tau}^{k}$
10:return $\hat{\mathbf{X}}_{t+\tau}^{k}$
Privacy enhancement: The privacy is ensured by LRSUQ executed cooperatively between the clients and the server in compliance with the threat model and the defined privacy guarantee. Through the decoding process, it is ensured that $\{\mathbf{Y}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ comprises noisy estimates of $\{\tilde{\mathbf{X}}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$, i.e., the estimated model updates at the server are noisy estimates of the clipped model updates at the clients, thus establishing a privacy mechanism. Refer to Step 7 in Algorithm 3. Specifically, in Section V-A1, we prove that when using the Gaussian mechanism, the global average of estimated model updates satisfies $(\epsilon,\delta)$-DP requirement by specifying appropriate $f$ and $g$, while in Section V-A2, when using the Laplace mechanism, the estimated model updates satisfies $\epsilon$-DP requirement by specifying appropriate $f$ and $g$.
V Performance Analysis
In this section, we study the performance of CEPAM. This includes characterizing its privacy guarantees and compression capabilities, followed by an exploration of its distortion bounds and convergence analysis. We begin by stating a useful lemma for the subsequent analysis. The proof is given in Appendix A-A.
Lemma 11.
@@ V-A Privacy
V-A1 Gaussian mechanism
If $g$ follows a chi-squared distribution with $n+2$ degrees of freedom, i.e., $g\sim\chi_{n+2}^{2}$, then $f$ follows a Gaussian distribution. This guarantees that the global average of estimated model updates satisfies $(\epsilon,\delta)$-DP. The proof is given in Appendix A-B.
Theorem 12.
Set $\tilde{U}_{t+\tau,j}^{k}\sim g=\chi_{n+2}^{2}$ and $\tilde{\mathbf{Z}}_{t+\tau,j}^{k}|\{\tilde{U}_{t+\tau,j}^{k}=u\}\sim\mathrm{Unif}(\sigma\sqrt{u}B^{n})$ for every $k,j$, the resulting mechanism, i.e., Step 7 in Algorithm 3, is Gaussian. The average $\frac{1}{K}\sum_{k\in\mathcal{K}}\hat{\mathbf{X}}_{t+\tau}^{k}$ is a noisy estimate of the average of the clipped model updates such that
$\frac{1}{K}\sum_{k\in\mathcal{K}}\hat{\mathbf{X}}_{t+\tau}^{k}=\frac{1}{K}\sum_{k\in\mathcal{K}}{\color[rgb]{0,0,0}\tilde{\mathbf{X}}_{t+\tau}^{k}}+{\color[rgb]{0,0,0}\mathcal{N}\bigg(\mathbf{0},\frac{\sigma^{2}}{K}\mathbf{I}_{m}\bigg)},$ (15)
@@ V-A1 Gaussian mechanism
From Lemma 9, the privacy profile $\delta$ of the subsampled Gaussian mechanism (Equation (15)) depends on the size $n$ of the dataset due to privacy amplification via subsampling with replacement, and on the global $\ell_{2}$-sensitivity $\Delta_{2}$. Note that the size of the dataset is $n=|\mathcal{D}^{(k)}|$ and the global $\ell_{2}$-sensitivity is at most $\frac{2\tau\gamma}{K}$.
V-A2 Laplace mechanism
If $g$ follows a Gamma distribution $\mathrm{Gamma}(2,1)$, then $f$ follows a Laplace distribution. [^8] Our CEPAM can be extended to include multivariate Laplace distributions [74] or $t$-distributions [75], which are left for future work. This guarantees that estimated model updates satisfy $\epsilon$-DP. The proof is given in Appendix A-C.
Theorem 13.
Set $\tilde{U}_{t+\tau,j}^{k}\sim g=\mathrm{Gamma}(2,1)$ and $\tilde{Z}_{t+\tau,j}^{k}|\{\tilde{U}_{t+\tau,j}^{k}=u\}\sim\mathrm{Unif}((-bu,bu))$ for every $k$ and $j$, the resulting mechanism is Laplace. The estimator $\hat{\mathbf{X}}_{t+\tau}^{k}$ is a noisy estimate of the clipped model update $\tilde{\mathbf{X}}_{t+\tau}^{k}$ such that
$\hat{\mathbf{X}}_{t+\tau}^{k}={\color[rgb]{0,0,0}\tilde{\mathbf{X}}_{t+\tau}^{k}}+{\color[rgb]{0,0,0}\mathrm{Lap}\left(\mathbf{0},b\mathbf{I}_{m}\right)},$ (17)
@@ V-B Compression
VI Numerical Evaluations
@@
In this section, we evaluate the performance of CEPAM. [^9] The source code used in our numerical evaluations is available at https://github.com/yokiwuuu/CEPAM.git We begin by detailing our experimental setup, which includes the types of datasets, model architectures, and training configurations in Section VI-A. Afterwards, we present comprehensive experimental results to demonstrate the effectiveness of CEPAM by comparing it to several baselines realizing the privacy protection and quantization in the FL framework in Section VI-B.
@@
We also compare the accuracy-privacy trade-off between CEPAM and the Gaussian-mechanism-then-quantize approach in Section VI-C. All experiments were executed on a server equipped with dual Intel Xeon Gold 6326 CPUs (48 cores and 96 threads in total), 256 GiB RAM, and two NVIDIA RTX A6000 GPUs (each with 48GB VRAM), running Ubuntu 22.04.5 LTS. The implementation was based on Python 3.13.5 and PyTorch 2.7.1+cu126, with CUDA 11.5 and NVIDIA driver version 565.57.01.
VI-A Experimental Setup
VI-A1 Datasets
@@ VI-A3 Baselines
  • FL+Gaussian: FL with one-dimensional Gaussian mechanism, no compression
  • FL+Gaussian+SDQ: FL with approach that applies one-dimensional Gaussian mechanism followed by scalar SDQ.
  • CEPAM-Gaussian: CEPAM achieves privacy and quantization jointly through LRSUQ to construct Gaussian mechanism, namely CEPAM-Gaussian. In particular, for CEPAM-Gaussian, we evaluate three different cases of LRSUQ that simulate Gaussian noise for dimension $n=1,2,3$, where we use a scaled integer lattice $\alpha\mathbb{Z}^{n}$, where $\alpha=10^{-5}$, [^10] To simplify implementation, we choose to use the integer lattice due to its well-established decoding algorithm [77]. The time complexity of the rejection sampling step can be improved by using a lattice with higher packing density. and the corresponding basic cells $\alpha(-0.5,0.5]$, $\alpha(-0.5,0.5]^{2}$, and $\alpha(-0.5,0.5]^{3}$. We have chosen $\alpha=10^{-5}$ as the precision level for quantization as it aligns with the observed precision level of the model updates during training.
VI-A4 Training Configurations